OpenStudy (anonymous):

How do you get the arc length of a circle with 10 or 12 triangles in the inside?

ganeshie8 (ganeshie8):

radius is given ? then may be you can use the formula : arc length = \((\theta/360) * r \)

OpenStudy (anonymous):

thank you

ganeshie8 (ganeshie8):

np...

ganeshie8 (ganeshie8):

oops! there is a mistake in the formula i have put

ganeshie8 (ganeshie8):

pls scratch my reply above arc length is just this : arc length = \(\theta * r\) i think you can find \(\theta\) seen by arc by dividing number of triangles in the circle : \(360/n\)

ganeshie8 (ganeshie8):